OpenStudy (anonymous):

The law of cosines is used to find the measures of a triangle when only the three sides are known, because the law of sines requires at least one angle measure. A. True B. False

OpenStudy (anonymous):

i think its false...

OpenStudy (anonymous):

I leaning towards true..

OpenStudy (anonymous):

i think its false becasue that is teh defintion for law of sinces not law of cosines

OpenStudy (anonymous):

The law of cosines is c^2 = a^2+ b^2-2ab cos(C), thus allowing you to find an angle of the triangle with only sides

OpenStudy (anonymous):

since the law of sines, a/sin A= b/sin B=c/sin C, you would have to know at least on angle to use it to solve for a variable

OpenStudy (anonymous):

oh